# Self-Supervised Learning in Event Sequences: A Comparative Study and Hybrid Approach of Generative Modeling and Contrastive Learning

### v1

@@ Abstract
This study investigates self-supervised learning techniques to obtain representations of Event Sequences (EvS). It is a key modality in various applications, including but not limited to banking, e-commerce, and healthcare.
We perform a comprehensive study of generative and contrastive approaches in self-supervised learning, applying them both independently. We find that there is no single supreme method. Consequently, we explore the potential benefits of combining these approaches. To achieve this goal, we introduce a novel method that aligns generative and contrastive embeddings as distinct modalities, drawing inspiration from contemporary multimodal research.
Generative and contrastive approaches are often treated as mutually exclusive, leaving a gap for their combined exploration. Our results demonstrate that this aligned model performs at least on par with, and mostly surpasses, existing methods and is more universal across a variety of tasks. Furthermore, we demonstrate that self-supervised methods consistently outperform the supervised approach on our datasets.
@@ 1 Introduction
Motivation. Event Sequences (EvS) are widely used in various real-world applications, including medicine Waring et al. (2020), biology Valeri et al. (2023), social medial analysis Liguori et al. (2023), fault diagnosis  Shao et al. (2019), churn prediction  Jain et al. (2021); Sudharsan and Ganesh (2022), customer segmentation  Carnein and Trautmann (2019), fraud detection  Xie et al. (2022) and more Zhuzhel et al. (2023); Fursov et al. (2021). As a result, there is a growing need to effectively model such data. Most self-supervised methods are focused on images, text, speech, or time-series. In Computer Vision (CV), contrastive learning methods have achieved state-of-the-art results as a pre-training strategy before supervised fine-tuning He et al. (2021); Chen et al. (2020); Caron et al. (2021). In Natural Language Processing (NLP), most modern models rely on generative pre-training Radford and Narasimhan (2018); Devlin et al. (2019). However, these approaches are largely unexplored for EvS.
The aim of our work is (1) to study generative and contrastive self-supervised approaches for pre-training and representation learning in the domain of EvS, (2) and to determine if they can complement each other. We study two ways to combine these approaches. For the first, we use a combined loss function. For the second, drawing inspiration from contemporary multimodal studies Radford et al. (2021); Zhai et al. (2023); Moskvoretskii and Kuznetsov (2023) we consider differently pre-trained models as distinct modalities and employ alignment techniques.
Contributions.
We are the first to our knowledge to examine the generative modeling for pre-training in the domain of EvS. We introduce a novel method called the Multimodal-Learning Event Model (MLEM) that aligns two pre-training strategies. Specifically, we use a pre-trained contrastive model to align a generative model. Our results have demonstrated that, on average, the aligned model outperforms any single method across a diverse range of tasks and datasets. Furthermore, our results indicate that self-supervised methods for pre-training outperform supervised approaches on all the datasets we have examined.
Furthermore, our study uncovers two significant insights. First, we find that the generative approach consistently achieves superior performance on tasks related to the Temporal Point Process (TPP), such as predicting the next event type and event time.
Second, we find that most of the methods are robust to perturbations along the time axis. However, they significantly degrade in performance on a synthetic dataset where the time component plays a crucial role. This observation suggests that some event sequences can be sufficiently classified using a ”Bag of Words” approach, where the order of events is not as important.
@@
We provide the source code for all the experiments described in this paper. [^1] The source code of our work is publicly available at https://github.com/VityaVitalich/MLEM.
2 Related work
Generative methods for pre-training and representation learning have seen notable development in recent years in both NLP Radford and Narasimhan (2018); Devlin et al. (2019); Raffel et al. (2023) and CV He et al. (2021); Assran et al. (2023); Bachmann et al. (2022). In NLP one of the generative pre-training procedures is the next token prediction used in BERT Devlin et al. (2019).
In Lin et al. (2022) authors study generative modeling of EvS for TPP tasks. Similarly to NLP, they use next-event prediction as their main training objective. While EvS has been studied in such a generative context, the authors in Lin et al. (2022) did not specifically investigate generative modeling as a representation of learning or pre-training. The similarities between autoregressive modeling in TPP and the success of generative approaches in NLP have prompted us to consider the potential for applying generative-style self-supervised learning and pre-training to the EvS domain.
Contrastive methods are a widely recognized in He et al. (2020); Grill et al. (2020). These methods typically draw closer the embeddings of variations of the same object and distance them from those of different objects. In CoLES Babaev et al. (2022) authors study contrastive approach for EvS. In their work, the positive samples consist of subsequences derived from a single user’s events, while the negative samples are obtained from events of different users. We utilize this method to examine the efficacy of the contrastive approach in our study.
Hybrid methods. Our work falls into the category of hybrid self-supervised approaches Qi et al. (2023); Yang et al. (2023); Lazarow et al. (2017); Oquab et al. (2023); Zhou et al. (2022); Kim et al. (2021); Kim and Ye (2022); Shukla and Marlin (2021). Few studies have merged generative and contrastive methods, typically focusing on loss combination. DINOv2 Oquab et al. (2023) updates the traditional contrastive loss from DINO Caron et al. (2021) with iBOT’s Zhou et al. (2022) reconstruction loss through Masked Image Modeling. Similarly, GCRL Kim et al. (2021) combines these losses via a weighted sum, significantly outperforming models that are purely generative or contrastive. EBCLR Kim and Ye (2022) also adopts this approach in the context of energy-based models, using a sum of contrastive and generative losses.
Other hybrid methods focus on combining generative and supervised learning Liu and Abbeel (2020). One such method, mTAND Shukla and Marlin (2021), has been applied for EvS classification.
Supervised learning is commonly used for the EvS classification task. Some works focus on modeling the underlying dynamics of the EvS process. For example, irregular samples can be modeled by employing dynamic modeling with Ordinary Differential Equations Rubanova et al. (2019). In contrast, SeqNAS Udovichenko et al. (2024) adopts Neural Architecture Search to identify optimal architectures without strong prior assumptions about the underlying process. A key finding is the effectiveness of Recurrent Neural Networks (RNNs) in modeling EvS, leading us to incorporate an RNN encoder in our model. Furthermore, the authors propose benchmark datasets for EvS classification. We selected the largest datasets from the benchmark, as self-supervised methods require substantial data volumes.
Figure 1: A scheme for contrastive pre-training and representation learning is illustrated using two different sequences,{$S_{1},S_{2}$}, where $S\in\mathcal{R}^{l\times f}$ and $l$ is a sequence and sequence length, respectively. The sub-sequences {$S_{1}^{{}^{\prime}},S_{2}^{{}^{\prime}},S_{1}^{{}^{\prime\prime}},S_{2}^{{}^{\prime\prime}}$} are sampled using a subsequence sampler described in Section 4.1, and the latent representations {$h_{1}^{{}^{\prime}},h_{2}^{{}^{\prime}},h_{1}^{{}^{\prime\prime}},h_{2}^{{}^{\prime\prime}}$}, where $h\in\mathcal{R}^{m}$, are produced by the contrastive encoder model $\mathcal{E}_{c}$. In our work, we utilize the GRU (Gated Recurrent Unit) and extract the last hidden state as $h$. The term $L^{con}$ denotes the contrastive loss.
Figure 2: In our generative method, the bottleneck encoder $\mathcal{E}_{g}$ encodes the entire sequence $S\in\mathcal{R}^{l\times f}$ into $h\in\mathcal{R}^{m}$, $l$ denotes sequence lengths and $f$ and $m$ corresponding hidden sizes. In our work, we utilize the GRU (Gated Recurrent Unit) and extract the last hidden state as $h$. For generation, we employ a transformer decoder with recursive generation conditioned on $h$. The term $L^{LM}$ denotes the reconstruction loss, and $[BOS]$ represents the Beginning Of Sequence token.
Figure 3: MLEM approach for EvS. We illustrate it with two different sequences {$S_{1},S_{2}$}. Both the contrastive CON and generative GEN encoders map the sequences {$S_{1},S_{2}$} into corresponding {$h_{1}^{c},h_{2}^{c},h_{1}^{g},h_{2}^{g}$} latent representations. We align the representations from different models using SigLIP and compute the alignment loss, $L^{align}$ to train GEN. Simultaneously we train GEN encoder with another reconstruction objective $L^{LM}$. In the end, only the GEN encoder is used as the final model for fine-tuning and obtaining representations. Throughout the procedure, we use a pre-trained CON encoder with frozen weights.
@@ 3 Event Sequences Preliminaries
Datasets used in this work can be described as sets of sequences: $C=\{(S_{1},y_{1}),(S_{2},y_{2}),\ldots\}$, where $y_{i}$ is an attribute corresponding to the entire sequence or some target. Each $S_{i}=((t^{1}_{i},\Upsilon^{1}_{i}),(t^{2}_{i},\Upsilon^{2}_{i}),\ldots)$ is a sequence of events $x^{j}_{i}=(t^{j}_{i},\Upsilon^{j}_{i})$, where $t_{i}^{j}$ represents the time when the event $x_{i}^{j}$ occurred and $t_{i}^{j}\leq t^{j+1}_{i}$. The set of sub-events $\Upsilon^{j}_{i}=\{k^{1},k^{2},\ldots\}$ describes the event $x_{i}^{j}$. It is important to note that in this work, the terms Event Sequences (EvS) and Irregularly Sampled Time-series (ISTS) are used interchangeably, as the occurrence of measurement or sampling for the latter can be seen as an event.
@@ 4.1 Contrastive Learning
Contrastive Learning encodes a sequence $S$ into a compact representation $f_{e}:S_{i}\rightarrow h_{i}\in\mathcal{R}^{m}$, by bringing positive pairs (i.e., semantically similar objects) closer to each other in the embedding space, while pushing negative pairs (i.e., dissimilar objects) further apart. In the CoLES Babaev et al. (2022) framework, the authors suggest constructing a set of positive pairs by sampling sub-sequences from a sequence $S_{i}\rightarrow\{S_{i}^{{}^{\prime}},S_{i}^{{}^{\prime\prime}}\}$, where each element in $\{S_{i}^{{}^{\prime}},S_{i}^{{}^{\prime\prime}}\}$ is shorter than $S_{i}$ and elements in $\{S_{i}^{{}^{\prime}},S_{i}^{{}^{\prime\prime}}\}$ may intersect, the number of sub-sequences is not limited to two. We adopt this approach and refer to it as the subsequence sampler in Figure 1. Further, we utilize loss function (1) from Hadsell et al. (2006). The overall procedure is illustrated in Figure 1.
$\begin{split}L^{con}&=\frac{1}{|C|}\sum_{i=1}^{|C|}\sum_{j=1}^{|C|}\Bigg{(}z_{ij}\frac{1}{2}||h_{i}-h_{j}||_{2}^{2}\\ &+(1-z_{ij})\frac{1}{2}\max\{0,\rho-||h_{i}-h_{j}||_{2}\}^{2}\Bigg{)}\end{split}$ (1)
Here, $z_{ij}\in\{0,1\}$ denotes if objects are semantically similar, $h$ is the sequence embedding, $\rho$ the minimal margin between dissimilar objects and $C=\{S_{1},S_{2},\ldots\}$ is a set of sequences.
@@ 4.2 Generative Modeling
We employ generative modeling to train our bottleneck encoder. Although our methodology is similar to Sequence-to-Sequence models, it differs as we use a bottleneck encoder to derive a single vector representation for the entire sequence.
For the encoder, we utilize RNN, which extracts the final hidden state $h$, which is subsequently passed into a Transformer decoder Vaswani et al. (2017). The decoder reconstructs the entire sequence, conditioned on $h$ via the Cross-Attention mechanism, in an auto-regressive manner. This entire process is illustrated in Figure 2.
To train the model we use the next event prediction objective, which is similar to training language models. Nonetheless, as each event in the sequence $x^{j}_{i}=(t^{j}_{i},\Upsilon^{j}_{i})$ contain multiple sub-events $\Upsilon=\{k^{1},k^{2},\ldots\}$, we need to reconstruct all of them. To this end, we apply the cross-entropy loss for each categorical $k$ and the Mean Squared Error (MSE) loss for each real-valued $k$. The final loss is a cumulative sum of the losses for all elements in $\Upsilon=\{k^{1},k^{2},\ldots\}$, plus the MSE loss for time $t$. We denote this loss as $L^{LM}$.
Intuitively, this procedure requires our encoder to develop a representation informative enough for the decoder to accurately reconstruct the entire sequence. Since we map all our sequences $S$ into $h\in\mathcal{R}^{m}$ we call all encoders bottleneck encoders.
All details related to the model can be found in Section 5.3.
@@
Table 1: Evaluation of self-supervised methods fine-tuned using supervised loss. ABank, PhysioNet and TaoBao are reported with ROC-AUC, AGE reported with accuracy and Pendulum with MSE.
4.3 Aligning generative encoder
To align the embeddings obtained through generative modeling with those acquired through contrastive learning, we treat them as distinct modalities and employ a contrastive aligning procedure inspired by CLIP Radford et al. (2021). We utilize the recent SigLIP loss Zhai et al. (2023) for this purpose. The resulting aligned encoder model is referred to as the Multimodal-Learning Event Model (MLEM).
Overall MLEM training procedure is exactly the same as for the generative model described in Section 4.2, except the alignment, which incurs an additional loss (2) resulting in the total loss (3). Training MLEM requires a pre-trained contrastive encoder. Using this encoder we train a generative model and align its embeddings with the embeddings produced by the contrastive encoder. Contrastive encoder weights are not updated during training. Both generative $\mathcal{E}_{g}$ and contrastive $\mathcal{E}_{c}$ encoders receive a set of sequences $C=\{S_{1},S_{2},\ldots\}$ and map each sequence $S$ into corresponding hidden states $h^{g}$ and $h^{c}$, then the goal of MLEM alignment is to pull embeddings obtained from the same sequence closer to each other and to push embeddings from different sequences further away. This procedure is illustrated in Figure 3. Below is the alignment loss we use:
@@ 4.3 Aligning generative encoder
All technical details can be found in Section 5.3.
4.4 Naïve method
As a baseline, we also examine a straightforward, or Naïve, approach that merges generative and contrastive methods. This is achieved by incorporating a contrastive objective into the generative model, akin to methodologies used in prior studies Kim et al. (2021); Oquab et al. (2023). The final loss is a weighted sum of objectives $L=\alpha L^{LM}+\beta L^{con}$. One of the downsides of this implementation is that we can not utilize full-length sequences because sub-sequence sampling is required for contrastive procedure. Further, we compare Naïve method with MLEM and show that while it is very close by design it leads to inferior results.
@@ 5.1 Datasets
In our study, we use five EvS datasets formally outlined in Section 3.
  • A subset of EvS datasets from SeqNAS benchmark Udovichenko et al. (2024), more precisely: ABank, Age, TaoBao. ABank and Age are bank transactions and TaoBao is user activity.
  • PhysioNet 2012 comprises highly sparse medical event data, with the goal of predicting mortality.
@@
Each dataset contains a range of features, both numerical and categorical. All datasets, with the exception of ABank, consist of irregularly sampled time steps. Comprehensive details for each dataset are available in the supplementary materials.
@@
Table 2: Evaluation of self-supervised methods using linear and non-linear probing for downstream tasks and TPP tasks. The table reports the mean and standard deviation calculated from three different seeds for downstream task metrics. For TPP tasks, only the mean values from three seeds are presented, as the standard deviation was consistently $0.00$. In the table, the best-performing values are highlighted in bold, while the second-best values are underlined.

### v3

# MLEM: Generative and Contrastive Learning as Distinct Modalities for Event Sequences

@@ Abstract
This study explores the application of self-supervised learning techniques for event sequences. It is a key modality in various applications such as banking, e-commerce, and healthcare. However, there is limited research on self-supervised learning for event sequences, and methods from other domains like images, texts, and speech may not easily transfer. To determine the most suitable approach, we conduct a detailed comparative analysis of previously identified best-performing methods. We find that neither the contrastive nor generative method is superior. Our assessment includes classifying event sequences, predicting the next event, and evaluating embedding quality. These results further highlight the potential benefits of combining both methods. Given the lack of research on hybrid models in this domain, we initially adapt the baseline model from another domain. However, upon observing its underperformance, we develop a novel method called the Multimodal-Learning Event Model (MLEM). MLEM treats contrastive learning and generative modeling as distinct yet complementary modalities, aligning their embeddings. The results of our study demonstrate that combining contrastive and generative approaches into one procedure with MLEM achieves superior performance across multiple metrics.
**footnotetext: These authors contributed equally to this work
@@ 1 Introduction
Motivation. Event Sequences (EvS) are widely used in various real-world applications, including medicine Waring et al. (2020), biology Valeri et al. (2023), social medial analysis Liguori et al. (2023), fault diagnosis  Shao et al. (2019), churn prediction  Jain et al. (2021); Sudharsan and Ganesh (2022), customer segmentation  Carnein and Trautmann (2019), fraud detection  Xie et al. (2022) and more Zhuzhel et al. (2023); Fursov et al. (2021). As a result, there is a growing need to effectively model such data. Most self-supervised methods are focused on images, text, speech, or time-series. In Computer Vision (CV), contrastive learning methods have achieved state-of-the-art results as a pre-training strategy before supervised fine-tuning He et al. (2021); Chen et al. (2020); Caron et al. (2021). In Natural Language Processing (NLP), most modern models rely on generative pre-training Radford and Narasimhan (2018); Devlin et al. (2019).
Self-Supervised Learning (approaches have not been extensively explored for EvS. However, applying existing SSL methods directly to EvS may be unreasonable or even impossible due to several key differences between EvS and other well-studied domains: a. Multiple Features, events in EvS can be described by multiple features, which adds complexity to the learning process. b. Categorical and Numerical Values, event features in EvS may include both categorical and numerical values, further complicating modeling. c. Non-Uniform Temporal Spacing, events in EvS are not uniformly spaced in time, which requires models to handle irregular temporal patterns. d. Temporal Proximity, it is to assumed that events occurring in close temporal proximity share relevant information. Given these unique characteristics of EvS, further research is needed to explore SSL approaches specifically tailored to ES and address these challenges. We describe a formal structure of EvS in Section 2.
The aim of our work is to answer the following questions: Q1: Which self-supervised approach is superior for ES modeling: generative or contrastive? Q2: Can we efficiently combine generative and contrastive approaches into one procedure to obtain more robust results?
The Contributions. To answer these questions, we (1) study established generative, contrastive, and hybrid approaches. (2) To the best of our knowledge, we are the first to examine generative pre-training, for EvS classification, and hybrid methods in the domain of EvS in general. (3) We introduce a new technique named MLEM, inspired by contemporary multimodal studies, such as (Radford et al., 2021; Zhai et al., 2023; Moskvoretskii and Kuznetsov, 2023). MLEM allows for the combination of various SSL approaches by treating differently pre-trained models as distinct modalities. More precisely, MLEM aligns the contrastive and generative embeddings. We demonstrate that such combination achieves superior performance on average across multiple datasets and metrics compared to existing methods.
Evaluation employs a diverse set of datasets to cover various aspects of the problem space: three real-world transactional datasets from the SeqNAS benchmark (Udovichenko et al., 2024), the widely-used medical dataset Physionet 2012, and a simulated dataset of pendulum motion designed to examine explicit time dependencies. This selection ensures that our findings are robust across different types of data and tasks. To comprehensively measure performance we employ a set of metrics, including downstream tasks of regression or classification of sequences, prediction of the next event type or timing, intrinsic dimension, and anisotropy.
Baselines. To ensure a clear and fair comparison, we have selected only the top-performing model from each setting for our analysis, comparison of baseline approaches is presented in Table 3 and Table 3. These include models from recent studies on contrastive learning (Babaev et al., 2022), generative modeling in EvS (Lin et al., 2022), and hybrid approaches from other domains Kim et al. (2021). We focus on these models because they are clearly superior in their categories, making the inclusion of lesser models unnecessary.
Figure 1: An example of one sequence $S_{i}$ with some categorical and real-valued sub-events $\{k^{1},k^{2},\ldots\}$.
@@
We provide the source code for all the experiments described in this paper.***The source code of our work is publicly available at https://github.com/VityaVitalich/MLEM.
@@ 2 Event Sequences Preliminaries
Datasets used in this work can be described as sets of sequences: $C=\{(S_{1},y_{1}),(S_{2},y_{2}),\ldots\}$, where $y_{i}$ is an attribute corresponding to the entire sequence or some target. Each $S_{i}=((t^{1}_{i},\Upsilon^{1}_{i}),(t^{2}_{i},\Upsilon^{2}_{i}),\ldots)$ is a sequence of events $x^{j}_{i}=(t^{j}_{i},\Upsilon^{j}_{i})$, where $t_{i}^{j}$ represents the time when the event $x_{i}^{j}$ occurred and $t_{i}^{j}\leq t^{j+1}_{i}$. The set of sub-events $\Upsilon^{j}_{i}=\{k^{1},k^{2},\ldots\}$ describes the event $x_{i}^{j}$. It is important to note that in this work, the terms Event Sequences (EvS) and Irregularly Sampled Time-series (ISTS) are used interchangeably, as the occurrence of measurement or sampling for the latter can be seen as an event. Figure 1 depicts an example of one sequence.
Figure 2: Schematic Models Overview: A) The Contrastive Model divides sequence $S_{i}$ into subsequences $\{S_{i}^{\prime},S_{i}^{\prime\prime}\}$, encodes them into latent vectors $\{h_{i}^{\prime},h_{i}^{\prime\prime}\}$, and applies contrastive loss $L_{\text{con}}$. B) The Generative Model autoregressively reconstructs sequence $S$ cross-attended to latent vector $h$, derived from $S$ via a bottleneck encoder. Language model loss $L_{\text{LM}}$ is applied to reconstructed sequence $\hat{S}$. C) The MLEM Hybrid Model integrates generative and contrastive approaches, treating them as distinct modalities. It generates latent vectors $h_{i}^{g}$ and $h_{i}^{c}$ from a learned generative encoder and frozen, pretrained contrastive encoder respectively. An alignment loss $L_{\text{align}}$ is utilized to align vectors from the same sequence while separating those from different sequences. Language modeling loss $L_{\text{LM}}$ is applied as in generative model.
@@ 3.1 Contrastive Learning
For our contrastive learning approach, we employ COLES Babaev et al. (2022) framework, as it has shown superior performance over other contrastive methods. It learns to encode the sequence $S$ into a latent vector $h\in\mathcal{R}^{m}$ by bringing sub-sequences of the same sequence closer in the embedding space, while pushing sub-sequences from different sequences further apart. The overall procedure is illustrated in Figure 2 and formal definition is given in Appendix Contrastive learning
@@ 3.2 Generative Modeling
For our generative model, we consider the best performing setting from GNTPP (Lin et al., 2022) and adapt it to our task.
For the encoder, we utilize RNN, as it shows dominance to our task in recent studies (Udovichenko et al., 2024). It extracts the final hidden state $h$, which is subsequently passed into a Transformer decoder (Vaswani et al., 2017). The decoder autoregressively reconstructs the entire sequence, conditioned on $h$ via the Cross-Attention mechanism. This entire process is illustrated in Figure 2.
To train the model, we use the next event prediction objective, which is similar to training language models. Nonetheless, as each event in the sequence $x^{j}_{i}=(t^{j}_{i},\Upsilon^{j}_{i})$ contain multiple sub-events $\Upsilon=\{k^{1},k^{2},\ldots\}$, we need to reconstruct all of them. To this end, we apply the cross-entropy loss for each categorical $k$ and the Mean Squared Error (MSE) loss for each real-valued $k$. The final loss is a cumulative sum of the losses for all elements in $\Upsilon=\{k^{1},k^{2},\ldots\}$, plus the MSE loss for time $t$. We denote this loss as $L^{LM}$.
Intuitively, this procedure requires our encoder to develop a representation informative enough for the decoder to accurately reconstruct the entire sequence. Since we map all our sequences $S$ into, $h\in\mathcal{R}^{m}$ we call all encoders bottleneck encoders.
Table 1: EvS modeling and choice of baselines
Supervised Contrastive Generative
Paper SeqNAS CoLES Generative TPP
2024 2022 2022
Best base models GRU GRU REV-ATT, LSTM
Table 2: Hybrid SSL approaches
Sum of Losses (Naïve) Multimodal
Paper DINO 2023,
iBOT 2022 MLEM
GCRL 2021
Domain Images and Text EvS
Table 3: Evaluation of methods fine-tuned with supervised loss for downstream task.
@@
3.3 Naïve Hybrid Method
We consider weighted summation of contrastive and generative loss as baseline hybrid model, as it showed dominant performance in previous studies (Kim et al., 2021). We call this approach Naïve since it is a straightforward combination of losses $L=\alpha L^{LM}+\beta L^{con}$.
3.4 MLEM
In our method, named MLEM, we propose to combine contrastive learning and generative modeling by treating them as distinct modalities with further alignment. To align the embeddings obtained through generative modeling with those acquired through contrastive learning, we employ a contrastive aligning procedure inspired by CLIP Radford et al. (2021). We utilize the recent SigLIP loss Zhai et al. (2023) for this purpose.
Overall MLEM training procedure is nearly the same as for the generative model described in Section 3.2, except the alignment, which incurs an additional loss (1) resulting in the total loss (2). Unlike CLIP and other multimodal models, our approach involves pre-training only one modality. Specifically, we pre-train the contrastive encoder first and then train the generative encoder, aligning it with the fixed contrastive encoder.
Formally, both generative $\mathcal{E}_{g}$ and contrastive $\mathcal{E}_{c}$ encoders receive a set of sequences $C=\{S_{1},S_{2},\ldots\}$ and map each sequence $S$ into corresponding latent vectors $h^{g}$ and $h^{c}$. The goal of MLEM alignment is to pull embeddings obtained from the same sequence closer to each other and to push embeddings from different sequences further away. This procedure is illustrated in Figure 2. Below is the alignment loss we use:
@@ 4.1 Datasets
In our study, we use five established EvS datasets formally outlined in Section 2.
  • A subset of EvS datasets from SeqNAS benchmark Udovichenko et al. (2024), more precisely: ABank, Age, TaoBao. ABank and Age are bank transactions, and TaoBao is user activity.
  • PhysioNet 2012 comprises highly sparse medical event data, with the goal of predicting mortality Silva et al. (2012).
@@
The detailed description is provided in Appendix Detailed Dataset Information.
Table 4: Evaluation of self-supervised methods using linear and non-linear probing for downstream and TPP tasks. The best-performing values are highlighted in bold, while the second-best values are underlined.